OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 5y = x − 7 4x − y = -10 a. (-8, 3) b. (1/5, 8) c. (-3, -2) d. (-1, -1)

OpenStudy (anonymous):

Hello :) u want answer or learn Modom ?

OpenStudy (anonymous):

I can try and learn it, I'm really bad at algebra :(

OpenStudy (anonymous):

But we can answer together :) OK ?

OpenStudy (anonymous):

Yes :D

OpenStudy (anonymous):

To find x or y we should be find one of them first :) Which one do u like ? X or y ?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

To find x we should killing Y s !!! Its so fun :)! To killing them y1+y2=0 OK ?!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now we have -y and 5y . So cant kill them :) OK ?

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

To killing them we should have -5y +5y Means -y*5 OK ?!

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

Did u understand ?:)

OpenStudy (anonymous):

Kind off :( With the info i kind of learned im gonna say its b :/

OpenStudy (anonymous):

Maenns yes or No ? It was important :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so we killed the y's and we have -y*5 rite

OpenStudy (anonymous):

Ok :) No problem :) We can start again :)

OpenStudy (anonymous):

@AsleyXoXo :So smart !

OpenStudy (anonymous):

:D thank u

OpenStudy (anonymous):

So for each of the coordinates given, (aka a,b,c,etc) They are all in the form of (x,y). So just substitute in whatever x is given into the formula so 5y= x-7, let's try plugging in a.(-8, 3) so 5(3) b/c the form is (x,y) = (-8) -7 = 15 = -15 this is not equal. so It isn't a solution for 5y = x - 7

OpenStudy (anonymous):

Now look : First we wanted X s so we should kill the y because we cant caculate them together :) OK?

OpenStudy (anonymous):

You can do the same with b. and plug that into the same equation and if it doesn't work, go onto c. and d. for the solutions.

OpenStudy (anonymous):

@ittskathherine :I said to u dont give the annnnnnnssssssswer .....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@E.ali I was just showing her a simpler way of solving for it, I used a. as an example, so I didn't give away the whole answer.

OpenStudy (anonymous):

you guys are smart :/

OpenStudy (anonymous):

@AsleyXoXo:Did u understand ittskathherine s way ?????!!!!!!

OpenStudy (anonymous):

I can tell you an even simpler way. You know when you graph, you get the coordinates (x,y)? That's what's here. Whatever number that's in the first blank between the parenthesis, is the x. (_,y). And whatever number that's in the second blank between the parentheses is the y. (x, _) I put blank spaces in there so that you can plug in whatever number they give you.

OpenStudy (anonymous):

i need to go back to pre algebra, they put me in algebra, i never toke pre algebra

OpenStudy (anonymous):

So for b. it lists, (1/5, 8) so one-fifths, eight. Which means, 1/5 is x and 8 is y. because of (x,y)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and c. is (-3,-2) which means, -3 is x and -2 is y. Get it? Can you tell me which number in d. is x and which is y? So i can see if you understand.

OpenStudy (anonymous):

1/5=x , 8=y rite/

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what do we do with those numbers

OpenStudy (anonymous):

In this case, for d. (the numbers are the exact same, so it doesn't matter which is x and which is y) But you should get the idea. The left side is x, and the right side is y. You know that formula they give you? 5y= x -7. You plug in the numbers from a. (which we already did), b.,c.,and d.

OpenStudy (anonymous):

So where y is, you would plug in the y number (aka 8) and where x is, you plug in the x number, (aka 1/5)

OpenStudy (anonymous):

so then you would end up with an equation like this 5 (8) = (1/5) -7. They want you to solve the equation. so 5 times 8 is 40, 1/5 - 7 is equal to 7/35 - 35/35 so 40= -28/35 But however, these are not equal so therefore it is not the solution. Try it with C. and D.

OpenStudy (anonymous):

@ittskathherine , you should let @E.ali to show his work!

OpenStudy (anonymous):

so c would be x=3, y =-2

OpenStudy (anonymous):

@PFEH.1999 He isn't replying though and she didn't really understand his way...

OpenStudy (anonymous):

But yes, you are right @AsleyXoXo

OpenStudy (anonymous):

ok,show your work,she won't understand you if i interrupt you :)

OpenStudy (anonymous):

so we plug those numbers into 5y = x − 7

OpenStudy (anonymous):

Yes, plug it in. Since you identified y and x, you can switch those numbers where x and y are

OpenStudy (anonymous):

-2y = -3x

OpenStudy (anonymous):

Not exactly, because 5y = x-7 so 5(-2) = (-3) -7

OpenStudy (anonymous):

so then you would put -2 where Y is because that's how it is in its coordinates. same with x

OpenStudy (anonymous):

so we plug those numbers into 5y = x − 7

OpenStudy (anonymous):

so multiply the 2 y's and x's rite

OpenStudy (anonymous):

We don't multiply them with y and x. we replace them with x and y

OpenStudy (anonymous):

we end up with this so 5(-2) = (-3) -7

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

what do we do to those numbers?

OpenStudy (anonymous):

Since 5 is multiplying y, (and we replaced y with -2, we multiply 5 with -2) and since x is subtracting 7, (and we replaced x with -3, we subtract 7 from -3.)

OpenStudy (anonymous):

-10, -10

OpenStudy (anonymous):

So yup! since they're equal, C., is a solution to 5y = x-7

OpenStudy (anonymous):

Makes sense now?

OpenStudy (anonymous):

Yes it does :D , i learned something new thanks a lot every one for helping

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

i got 70% on my quiz, thank u all once again <3